OpenStudy (anonymous):

The equation of a line that goes through the points (-4,3) and (4,-4) can be written in the form y=mx+b, where m is? B is?

OpenStudy (anonymous):

m is slope, b is your constant

OpenStudy (lgbasallote):

solve it via two point form first then transform to y= mx + b the two point form is y- y1 = (y2 - y1) (x- x1) ------- (x2 - x1) where x1 and y1 are the first x and y coordinates respectively, and x2 and y2 are the second x and y coordinates respecitvely

OpenStudy (anonymous):

I can get that far, then I get lost

OpenStudy (lgbasallote):

what did you get so far?

OpenStudy (anonymous):

a fraction

OpenStudy (lgbasallote):

tell me the equation you got...

OpenStudy (lgbasallote):

how could i when you wont show me what you did....